Red blood cell count is distributed normally with a mean of 5.4 million cells /uL of blood and a standard deviation of 0.7 million cells /uL of blood. What is the probability of a healthy adult having more than 6.5 million cells /uL of blood? (Write your answer as 0.XXX)
Answer #1

= 5.4

= 0.7

To find P(X>6.5):

Z = (6.5 - 5.4)/0.7 = 1.5714

table of AreaUnder Standard Normal Curve gives area = 0.4418

So,

P(X>6.5) = 0.5 - 0.4418 = 0.0582

So,

Answer is:

0.058